OpenStudy (anonymous):

What is the distance between the point (-5, -4) and the line 2x + 3y + 5 = 0

OpenStudy (anonymous):

A. 23 √13 B. 17 √13 C. 17 23 D. √23 17

OpenStudy (anonymous):

option B

OpenStudy (anonymous):

perpendicular distance of point (x1,y1) from line ax+by+c=0 is given by \[|a*x1+b*y1+c| / \sqrt{a^{2}+b^{2}} \] Here (x1,y1)=(-5,-4) a=2 b=3 c=5

OpenStudy (anonymous):

DIstance between a point (x,y) and a line ax+by+c is\[\frac{ ax+by+c }{\sqrt{a^2+b^2} }\] modulus of it.

OpenStudy (anonymous):

since distance=|(-5*2)+(-4*3)+5| /\[\sqrt{3^{2}+2^{2}}\]